OpenStudy (anonymous):

start read Customer_Order_Record if Quantity > 100 then Discount = 20 else if Quantity > 12 then Discount = 10 endif endif Total = Price* Quantity Total = Total - (Discount* Total)/100 Print Total Stop >> Example 2 Start Read Student_Marks_File Pass = 0 Fail = 0 Total = 0 X = 0 while (End Of Student_Marks_File) do Total = Total + Subject_Marks if (Subject_Marks > 40) then Print "Student passed in" Subject else

OpenStudy (anonymous):

Print "Student failed in" Subject endif X = X +1 Read Next Record endwhile Average = Total/X Print "Average Marks of" Student_Name "are" Average Stop From the above examples of pseudocode containing control structures, analyze the logic implemented in the pseudocode.

OpenStudy (anonymous):

I need help with this please

OpenStudy (anonymous):

what is the problem , m unable to get which kind of help u need ?

OpenStudy (anonymous):

I need help to understand the logic behind each set of pseudocode

OpenStudy (anonymous):

so what do you need to know

OpenStudy (anonymous):

How would you implement the logic used for the above examples?

OpenStudy (anonymous):

There is 2 examples and I usually have a tendency to over think the question asked

OpenStudy (anonymous):

Analyze Pseudocode The following are a few examples of pseudocode: >> Example 1 start read Customer_Order_Record if Quantity > 100 then Discount = 20 else if Quantity > 12 then Discount = 10 endif endif Total = Price* Quantity Total = Total - (Discount* Total)/100 Print Total Stop >> Example 2 Start Read Student_Marks_File Pass = 0 Fail = 0 Total = 0 X = 0 while (End Of Student_Marks_File) do Total = Total + Subject_Marks if (Subject_Marks > 40) then Print "Student passed in" Subject else Print "Student failed in" Subject endif X = X +1 Read Next Record endwhile Average = Total/X Print "Average Marks of" Student_Name "are" Average Stop From the above examples of pseudocode containing control structures, analyze the logic implemented in the pseudocode. Answer Preview ...to the "Total" variable (overwriting its existing value). Continuing from the previous example, "Total" would be 100 - 20 = 80. 6. Finally, total is printed by the "Print Total" command. The above explanation is rather too detailed for such a...

OpenStudy (anonymous):

dose that help any

OpenStudy (anonymous):

It does if it could be viewed entirely sorry

OpenStudy (anonymous):

Would it help if it was broken down?